# On the reconstruction limits of complex networks

### v2

@@ II.4.1 Reconstruction from empirical neuronal spiking data
We consider the spontaneous activity of 1462 neurons from a mouse brain recorded over 20 minutes using eight neopixel probes [Fig. 6(a)] Stringer et al. (2019). Starting from the recorded spike times of the neurons, indicating when they fire, we create a binary time series of the activity of each neuron. In these binary time series, a ‘1’ marks the moments when a neuron is fired, and a ‘0’ when it is not. For more detail regarding our data processing procedure, see Appendix IV.9. Then, we infer the network using a variety of Bayesian reconstruction models. As a result, we get the reconstructed networks shown in Fig. 6(b), where we aggregated the edges connecting the neurons of all pairs of probes, thus illustrating how they interact. Finally, we compute the log evidence for each model and select the model with the highest one (see Appendix IV.6 for details on the evidence estimation). In doing so, we extract the reconstruction index that is, in principle, closest to the reconstruction limit, given the set of considered models. Figure 6(c) shows a diagram of both measures for each model.
@@
In our example, the analyzed reconstruction models are combinations of time series likelihoods—SIS, Cowan and Glauber—and graph priors—the ER model, the configuration model with a uniform degree sequence hyperprior (UCM) and the stochastic block model (SBM). The ER model, having a uniform distribution, is the most entropic model, followed by the SBM and the CM. Additional details about the graph prior and about these reconstruction models are given in Appendices IV.1 and IV.2, respectively. Of course, these models oversimplify the observed neuronal activity. Moreover, certain critical factors are not captured in the current dataset, such as the latency and deactivation rates of neuronal activity, as well as the substantial number of neurons undetected by the probes, which could contribute as input currents to the modeled neurons. In addition, the lack of detailed connectomic data for such small brain regions prevents a rigorous assessment of the accuracy reconstructed probe network. The purpose of this analysis is therefore not to perform the most accurate reconstruction but to illustrate the complete procedure as well as the results it generates.
@@
Among the considered models, the SIS model with a SBM prior is the one achieving the highest log evidence. As detailed in Appendix IV.9, the inferred SIS model reproduces many of the statistics featured in the data, suggesting that its corresponding reconstruction index of approximately $67\%$ is reasonable. Additionally, given that the network contains $1462$ neurons and that the estimated number of edges for this model is approximately $1722$ (see Table 3), the inferred network is sparser than expected (e.g., see Table 2 in Lin et al. (2024)) with an average degree of $2.35$. This suggests that, although our estimation of reconstructability is not close to zero, the inferred network does not seem to account for most of the neural activity. Of course, considering the decreasing tendency observed in Fig. 6(c), more detailed neuronal models—better suited for these data and with potentially higher log evidence—could yield reconstruction indices even lower than $67\%$, with possibly denser inferred networks. In the following section, we circumvent the limitations of the neuronal activity data by transitioning to a controlled setting, where synthetic activity data is used to reconstruct empirical networks.
II.4.2 Reconstruction of empirical graphs from synthetic activity data
@@
Table 1: Negative log probability—i.e., $-\log P(G=g^{*})$—of the graphs considered in Fig. 7 using the Erdős-Rényi (ER) model, the configuration model with uniform degree sequence prior (UCM), the configuration model with given degree sequence (CM) and the stochastic block model (SBM).
@@
Figures 7(a) and (f) show the reconstruction index of two empirical graphs, namely Zachary’ karate club Zachary (1977) and the Political books network Krebs (2004), as a function of the posterior loss for different graph prior models. Here, the reconstruction indices $\Psi_{M}$ are calculated using Eq. (29) like before, where all $x^{*}$ are generated using the same graph $g^{*}$. For this analysis, in addition to the previous graph priors, we consider the standard CM where the degree sequence is given. The standard CM is less entropic than the UCM since it is given all the information of the degree sequence—it does not need to be inferred from $X$ as for the UCM. The graph support of the CM is also quite a bit smaller than that of the other priors, which should signification improve the reconstruction. We generated many synthetic observations on these two graphs $g^{*}$ to perform the reconstruction: A spreading epidemics on the karate club social network using the SIS model, and the Voter model for the Political books network, as to simulate the propagation of political opinions. For both examples, we chose many values of parameters for the dynamics to present the complete range of reconstructability scenarios. To highlight the role of the graph prior, we reconstruct all graphs with the TDG data models. This means that, in this case, $\Psi_{M}$ is indeed a point-wise measure of reconstructability.
@@
The reconstruction limit, as measured by $\Psi_{M}$, changes as a function of the graph prior: If information about $g^{*}$ is a priori given, this same information cannot be reconstructed from realizations of $X^{*}$. Thus, it is no longer taken into account in $\Psi_{M}$ which by construction factors out the contributions of the prior. This intuition can be mathematically studied if we let the prior put more and more weight on $g^{*}$. In Appendix IV.8, we prove that as the graph generative model converges to a Kronecker delta distribution, i.e.,
@@
the reconstructability $U(G^{*}|X^{*})$ converges to zero, except if the mutual information is maximized in which case it is always equal to 1.
III Discussion
@@
We envision a future where network reconstruction applications incorporate a reconstructability analysis in their pipeline, such as that presented in this paper. By doing so, the reconstruction index would indicate how informative the reconstructed networks are and perhaps inform us on how they should be used within the said applications. Of course, there is still plenty of work to be done on this front, such as improving the computational methods required to compute the reconstruction index as they do not scale well to large networks, and improving the reconstruction models themselves as we have alluded to earlier. Some of these models might also require modifying our framework, for example in the case of weighted and directed networks. These specific models could prove considerably valuable for the neuroscience community and, more broadly, for complex systems research.
IV Materials and methods
IV.1 Graph priors
In the paper, we use different random graph models as graph priors for Bayesian network reconstruction. These models are undirected and unweighted and may include self-loops and multiedges, although our general framework is not restricted to these assumptions. Indeed, one could consider directed or weighted graphs as well; as long as the set of possible graphs remains countable. We use the adjacency matrix, denoted $\bm{a}$, in order to define the probability distribution of some of these models, where $a_{ij}$ counts the number of edges connecting nodes $i$ and $j$. To simplify the notation, we will sometimes express a graph $g$ directly with its adjacency matrix $g=\bm{a}$ We use the convention that $a_{ii}$ is always a multiple of 2. Below, we describe these priors in more detail.
IV.1.1 Erdős-Rényi model
@@
such that ((${}{n}{k}\right)\kern-3.00003pt\right)=\binom{n+k-1}{k}$ counts the number of possible multisets of size $k$ composed of $n$ different objects—i.e., multiedges.
Note that the number of edges $E$ must be provided to the ER model, and in the other graph models described below. This means that $\theta=E$ is the hyperparameter of ER graph prior and that $E$ should be inferred. We use the prior $P(E)$ to weigh in the number of edges. In most of our experiments, the number of edges is fixed to a specific value $e^{*}$, meaning that $P(E=e)=\delta(e,e^{*})$, where $\delta(m,n)$ is the Kronecker delta function. However, in Sec. II.4.1, as $E$ is unknown in this case, we use a geometric prior of the form
$P(E=e)=\frac{\bar{\lambda}^{e}}{(\bar{\lambda}+1)^{e+1}}\,$ (36)
where $\bar{\lambda}$ is a parameter that fixes the expected number of edges. See also Appendix IV.9 for further detail about the complete inference procedure.
IV.1.2 Configuration model
@@
$P(G=\bm{a}|\bm{k}=\bm{\kappa})=\frac{\prod_{i=1}^{N}\kappa_{i}!}{\prod_{i<j}a_{ij}!\prod_{i=1}^{N}a_{ii}!!}\,$ (38)
where $(2n)!!=2^{n}n!$.
@@
where ((${}{N}{2e}\right)\kern-3.00003pt\right)$ counts the number of possible degree sequences for a graph of $N$ nodes and $E$ edges.
IV.1.3 Stochastic block model
The stochastic block model (SBM), in its microcanonical version Peixoto (2017), closely resembles the ER model, where edges are picked uniformly at random. However, unlike the ER model, each node $i$ is associated with a random block $b_{i}\in\{1,2,\dots,B\}$ and instead the number of edges $e_{rs}=\sum_{ij}a_{ij}\delta(b_{i},r)\delta(b_{j},s)$ connecting two blocks $r$ and $s$ is fixed such that there are $E$ edges in total. We summarize the node partition as a tuple $\bm{b}=(b_{i})_{i=1..N}$ and the number of edges between blocks by the edge matrix $\bm{e}=(e_{rs})_{r,s=1..B}$. The blocks are required to be non-empty. Since $B$ is the number of non-empty blocks in $\bm{b}$ and $\bm{e}$ is completely determined by the graph and $\bm{b}$, $\theta=(\bm{b},E)$ are the hyperparameters of the SBM. Consequently, $\bm{b}$ and $E$ must be inferred jointly with the graph. In theory, one could factor the joint prior probability $P(G,E,\bm{b})$ as $P(G|E,\bm{b})P(E)P(\bm{b})$ assuming $E$ and $\bm{b}$ are independent. However, it is more convenient to factor the prior also using $\bm{e}$ and $B$ as intermediate random variables, following
@@ IV.1.3 Stochastic block model
IV.2 Markov chain likelihoods
@@
$\alpha(n,m,\varphi)$ $=(1-\alpha_{0})\tilde{\alpha}(n,m,\varphi)+\alpha_{0}$ (49)
@@
In general, we fix $\alpha_{0}=\beta_{0}=0$ for the synthetic experiments, and infer them in Sec. II.4.1.
@@
Table 2: Activation and deactivation probability functions for the likelihoods used in this paper, where $n$ corresponds to the number of inactive neighbors whose states are 00, and $m$ corresponds to the number of active neighbors whose states are $1$. We define $\sigma(x)=[\exp(-x)+1]^{-1}$ as the sigmoid function.
Table 2 presents the activation and deactivation probability functions for four different processes used in various contexts. The Glauber dynamics is a spin model that describe the time-reversible evolution of magnetic spins (00 or $1$) aligning in a crystal. In this model, the nodes are connected through their neighbors via a coupling constant $J$, that modulates the probability of a node to align with its neighbors. The susceptible-infected-susceptible (SIS) dynamics is a canonical model of epidemic spreading, where the nodes are either susceptible (00) or infected ($1$), and has often been used to model disease with short immunity after recovery, similar to influenza-like disease Anderson and May (1992). Susceptible (or inactive) nodes get infected by each of their infected (active) first neighbors, with a constant transmission probability, and recover from the disease with a constant recovery probability. The Voter dynamics model the adoption of opinions; A node randomly selects the opinion (two opinions, 00 or $1$, are considered) of one of its neighbors. The Cowan dynamics is a model of neural activity of biological neural networks, where the nodes—referred to as neurons—are either active ($1$) or inactive (00), and has been used to model the dynamics of single neurons or neuronal populations Cowan (1990); Painchaud et al. (2022). Inactive neurons fire—i.e., become active—if their input current, coming from their firing neighbors, is above a given threshold.
The parameters $\phi$ of these models are fixed except in Sec. II.4.1 where they are inferred by sampling from the joint posterior $P(G,\phi,\theta|X)$. In these experiments, we use non-informative uniform prior densities for all parameters in $\phi$, and we constrain their value in finite intervals. For probability parameters, such as $\alpha_{0}$ and $\beta$ for the SIS and Cowan models, the prior density is $\rho(\phi)=1$. For positive unbounded parameters, such as $J$ for the Glauber model, and $\mu$ and $\nu$ for the Cowan model, we set the maximum value to $10$ such that $\rho(\phi)=\frac{1}{10}$. Note that we fix $a=1$ in the case of the Cowan model and $\beta_{0}=0$ for the SIS and the Cowan models in Sec. II.4.1, without loss of generality since they are redundant parameters that may lead to non-identifiability issues.
IV.3 Heuristic reconstruction algorithms
@@
IV.4 Relationship between the posterior loss and the reconstructability
@@
We also write $L\big{(}\bm{y},\bm{p}^{*}\big{)}$ the posterior loss of the TDG model $M^{*}$, such that $\bm{p}^{*}$ is its corresponding posterior probability vector. Rewriting the posterior loss in terms of the posterior probability, simply get
@@
$\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p})]$ $=-\mathbb{E}_{X^{*},G^{*}}\quantity[\log P(G=G^{*}|X=X^{*})]\,.$ (58)
Hence, when $M$ and $M^{*}$ are equal in distribution, $\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p}^{*})]=H(G^{*}|X^{*})$, which is the posterior entropy of the TDG model. Furthermore, the expected posterior loss is linearly related to the reconstructability, with a proportionality factor given by the entropy of $G^{*}$:
$\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p}^{*})]=H(G^{*})\Big{[}1-U(G^{*}|X^{*})\Big{]}\,.$ (59)
IV.5 Bounds of the information gain
In this section, we show that the information gain is non-negative and bounded by the CE between the posterior and the prior. Recall that the information gain is given by Eq. (II.3.1) (equivalently Eq. (30)):
$\mathcal{I}_{M}(x)=\mathbb{E}_{G|X=x}\quantity[\log\frac{P(G|X)}{P(G)}]\,$

### v3

@@ V.1 Reconstruction from empirical neuronal spiking data
We consider the spontaneous activity of 1462 neurons from the dorsal cerebral cortex of a mouse recorded over 20 minutes using eight Neuropixel probes [Fig. 6(a)] Stringer et al. (2019). Starting from the recorded spike times of the neurons, indicating when they fire, we create a binary time series of the activity of each neuron. In these binary time series, a ‘1’ marks the moments when a neuron is fired, and a ‘0’ when it is not. For more detail regarding our data processing procedure, see Appendix I. Following our method described at the beginning of Sec. V, we infer the network using a variety of Bayesian reconstruction models. As a result, we get the reconstructed networks shown in Fig. 6(b), where we aggregated the edges connecting the neurons of all pairs of probes, thus illustrating how they interact. Finally, we compute the log evidence for each model and select the model with the highest one (see Appendix F for details on the evidence estimation). In doing so, we extract the reconstruction index that is, in principle, closest to the reconstruction limit, given the set of considered models. Figure 6(c) shows a diagram of both measures for each model.
@@
In our example, the analyzed reconstruction models are combinations of time series likelihoods—SIS, Cowan and Glauber—and graph priors—the ER model, the configuration model with a uniform degree sequence hyperprior (UCM) and the stochastic block model (SBM). The ER model, having a uniform distribution, is the most entropic model, followed by the SBM and the CM. Additional details about the graph prior and about these reconstruction models are given in Appendices A and B, respectively. Of course, these models oversimplify the observed neuronal activity. Moreover, certain critical factors are not captured in the current dataset, such as the latency and deactivation rates of neuronal activity, as well as the substantial number of neurons undetected by the probes, which could contribute as input currents to the modeled neurons. In addition, the lack of detailed connectomic data for such small brain regions prevents a rigorous assessment of the accuracy reconstructed probe network. The purpose of this analysis is therefore not to perform the most accurate reconstruction but to illustrate the complete procedure as well as the results it generates.
@@
Among the considered models, the SIS model with a SBM prior is the one achieving the highest log evidence. As detailed in Appendix I, the posterior predictive checks of the inferred SIS model suggests that its reconstruction index of approximately $67\%$ is reasonable. Additionally, given that the network contains $1462$ neurons and that the estimated number of edges for this model is approximately $1722$ (see Table 3), the average degree is $2.35$, indicating a network far sparser than anatomical data would predict. Indeed, a recent connectomic study of a subvolume of the mouse visual cortex Turner et al. (2022)—part of the dorsal cortex studied in Stringer et al. (2019)—reports a connection probability of $0.054$. If we hypothetically assume a similar probability across the entire dorsal cortex, the expected total number of edges would be approximately $1.15\times 10^{5}$, corresponding to an average degree of about $79$. This suggests that, although our estimation of reconstructability is not close to zero, the inferred network does not seem to account for most of the neural activity. Of course, considering the decreasing tendency observed in Fig. 6(c), more detailed neuronal models—better suited for these data and with potentially higher log evidence—could yield reconstruction indices even lower than $67\%$, with possibly denser inferred networks. In the following section, we circumvent the limitations of the neuronal activity data by transitioning to a controlled setting, where synthetic activity data is used to reconstruct empirical networks.
V.2 Reconstruction of empirical graphs from synthetic activity data
@@
Table 1: Negative log probability—i.e., $\log P(G=g^{*})$—of the graphs considered in Fig. 7 using the Erdős-Rényi (ER) model, the configuration model with uniform degree sequence prior (UCM), the configuration model with given degree sequence (CM) and the stochastic block model (SBM).
@@
Our method can also be used to assess the role of the prior, when using empirical graphs. However, since we consider here synthetic data generation process, we can omit the posterior predictive checks. Figures 7(a) and (f) show the reconstruction index of two empirical graphs, namely Zachary’ karate club Zachary (1977) and the Political books network Krebs (2004), as a function of the posterior loss for different graph prior models. Here, the reconstruction indices $\psi_{M}$ are calculated using Eq. (29) like before, where all $x^{*}$ are generated using the same graph $g^{*}$. For this analysis, in addition to the previous graph priors, we consider the standard CM where the degree sequence is given. The standard CM is less entropic than the UCM since it is given all the information of the degree sequence—it does not need to be inferred from $X$ as for the UCM. The graph support of the CM is also quite smaller than that of the other priors, which should improve the reconstruction. We generated many synthetic observations on these two graphs $g^{*}$ to perform the reconstruction: A spreading epidemics on the karate club social network using the SIS model, and the Voter model for the Political books network, as to simulate the propagation of political opinions. For both examples, we chose many values of parameters for the dynamics to present the complete range of reconstructability scenarios. To highlight the role of the graph prior, we reconstruct all graphs with the TDG data models. This means that, in this case, $\psi_{M}$ is indeed a point-wise measure of reconstructability.
@@
The reconstruction limit, as measured by $\psi_{M}$, changes as a function of the graph prior: If information about $g^{*}$ is a priori given, this same information cannot be reconstructed from realizations of $X^{*}$. Thus, it is no longer taken into account in $\psi_{M}$ which by construction factors out the contributions of the prior. This intuition can be mathematically studied if we let the prior put more and more weight on $g^{*}$. In Appendix H, we prove that as the graph generative model converges to a Kronecker delta distribution, i.e.,
@@
the reconstructability $\Psi^{*}$ converges to zero, except if the mutual information is maximized in which case it is always equal to 1.
VI Conclusion
@@
We envision a future where network reconstruction applications incorporate a reconstructability analysis in their pipeline, such as the one presented in Sec. V. By doing so, the reconstruction index would indicate how informative the reconstructed networks are and perhaps inform us on how they should be used within the said applications. Of course, there is still plenty of work to be done on this front, such as improving the computational methods required to compute the reconstruction index as they do not scale well to large networks, and improving the reconstruction models themselves as we have alluded to earlier. Some of these models might also require modifying our framework, for example in the case of weighted and directed networks. These specific models could prove considerably valuable for the neuroscience community and, more broadly, for complex systems research.
Appendix A Graph priors
In the paper, we use different random graph models as graph priors for Bayesian network reconstruction. These models are undirected and unweighted and may include self-loops and multiedges, although our general framework is not restricted to these assumptions. Indeed, one could consider directed or weighted graphs as well; as long as the set of possible graphs remains countable. We use the adjacency matrix, denoted $\bm{a}$, in order to define the probability distribution of some of these models, where $a_{ij}$ counts the number of edges connecting nodes $i$ and $j$. To simplify the notation, we will sometimes express a graph $g$ directly with its adjacency matrix $g=\bm{a}$. We use the convention that $a_{ii}$ is always a multiple of 2. Below, we describe these priors in more detail.
A.1 Erdős-Rényi model
@@
[…]
Note that the number of edges $E$ must be provided to the ER model, and in the other graph models described below. This means that $\theta=E$ is the hyperparameter of ER graph prior and that $E$ should be inferred. We use the prior $P(E)$ to weigh in the number of edges. In most of our experiments, the number of edges is fixed to a specific value $e^{*}$, meaning that $P(E=e)=\delta(e,e^{*})$, where $\delta(m,n)$ is the Kronecker delta function. However, in Sec. V.1, as $E$ is unknown in this case, we use a geometric prior of the form
$P(E=e)=\frac{\bar{\lambda}^{e}}{(\bar{\lambda}+1)^{e+1}}\,$, (37)
where $\bar{\lambda}$ is a parameter that fixes the expected number of edges. See also Appendix I for further detail about the complete inference procedure.
A.2 Configuration model
@@
$P(G=\bm{a}|\bm{k}=\bm{\kappa})=\frac{\prod_{i=1}^{N}\kappa_{i}!}{\prod_{i<j}a_{ij}!\prod_{i=1}^{N}a_{ii}!!}\,$, (39)
where $(2n)!!=2^{n}n!$ is the double factorial of $2n$, i.e., the product of all even numbers up to $2n$.
@@
[…]
A.3 Stochastic block model
The stochastic block model (SBM), in its microcanonical version Peixoto (2017), closely resembles the ER model, where edges are picked uniformly at random. However, unlike the ER model, each node $i$ is associated with a random block $b_{i}\in\{1,2,\dots,B\}$ and instead the number of edges $e_{rs}=\sum_{ij}a_{ij}\delta(b_{i},r)\delta(b_{j},s)$ connecting two blocks $r$ and $s$ is fixed such that there are $E$ edges in total. We summarize the node partition as a tuple $\bm{b}=(b_{i})_{i=1..N}$ and the number of edges between blocks by the edge matrix $\bm{e}=(e_{rs})_{r,s=1..B}$. The blocks are required to be non-empty. Since $B$ is the number of non-empty blocks in $\bm{b}$ and $\bm{e}$ are completely determined by the graph and $\bm{b}$, $\theta=(\bm{b},E)$ are the hyperparameters of the SBM, then $\bm{b}$ and $E$ must be inferred jointly with the graph. In theory, one could factor the joint prior probability $P(G,E,\bm{b})$ as $P(G|E,\bm{b})P(E)P(\bm{b})$ assuming $E$ and $\bm{b}$ are independent. However, it is more convenient to factor the prior also using $\bm{e}$ and $B$ as intermediate random variables, following
@@ A.3 Stochastic block model
Appendix B Markov chain likelihoods
@@
$\alpha(n,m,\varphi)$ $=(1-\alpha_{0})\tilde{\alpha}(n,m,\varphi)+\alpha_{0}\,$, (50)
@@
In general, we fix $\alpha_{0}=\beta_{0}=0$ for the synthetic experiments, and infer them in Sec. V.1.
@@
Table 2: Activation and deactivation probability functions for the likelihoods used in this paper, where $n$ corresponds to the number of inactive neighbors whose states are 0, and $m$ corresponds to the number of active neighbors whose states are $1$. We define $\sigma(x)=[\exp(-x)+1]^{-1}$ as the sigmoid function.
Table 2 presents the activation and deactivation probability functions for four different processes used in various contexts. The Glauber dynamics is a spin model that describe the time-reversible evolution of magnetic spins (0 or $1$) aligning in a crystal. In this model, the nodes are connected through their neighbors via a coupling constant $J$, that modulates the probability of a node to align with its neighbors. The susceptible-infected-susceptible (SIS) dynamics is a canonical model of epidemic spreading, where the nodes are either susceptible (0) or infected ($1$), and has often been used to model disease with short immunity after recovery, similar to influenza-like disease Anderson and May (1992). Susceptible (or inactive) nodes get infected by each of their infected (active) first neighbors, with a constant transmission probability, and recover from the disease with a constant recovery probability. The Voter dynamics model the adoption of opinions; A node randomly selects the opinion (two opinions, 0 or $1$, are considered) of one of its neighbors. The Cowan dynamics is a model of neural activity of biological neural networks, where the nodes—referred to as neurons—are either active ($1$) or inactive (0), and has been used to model the dynamics of single neurons or neuronal populations Cowan (1990); Painchaud et al. (2022). Inactive neurons fire—i.e., become active—if their input current, coming from their firing neighbors, is above a given threshold.
The parameters $\phi$ of these models are fixed except in Sec. V.1 where they are inferred by sampling from the joint posterior $P(G,\phi,\theta|X)$. In these experiments, we use non-informative uniform prior densities for all parameters in $\phi$, and we constrain their value in finite intervals. For probability parameters, such as $\alpha_{0}$ and $\beta$ for the SIS and Cowan models, the prior density is $\rho(\phi)=1$. For positive unbounded parameters, such as $J$ for the Glauber model, and $\mu$ and $\nu$ for the Cowan model, we set the maximum value to $10$ such that $\rho(\phi)=\frac{1}{10}$. Note that we fix $a=1$ in the case of the Cowan model and $\beta_{0}=0$ for the SIS and the Cowan models in Sec. V.1, without loss of generality since they are redundant parameters that may lead to non-identifiability issues.
Appendix C Heuristic reconstruction algorithms
@@
Appendix D Relationship between the posterior loss and the reconstructability
@@
We also write $L\big{(}\bm{y},\bm{p}^{*}\big{)}$ the posterior loss of the TDG model $M^{*}$, such that $\bm{p}^{*}$ is its corresponding posterior probability vector. Rewriting the posterior loss in terms of the posterior probability, we simply get
@@
$\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p})]$ $\approx-\mathbb{E}_{X^{*},G^{*}}\quantity[\log P(G=G^{*}|X=X^{*})]\,$, (59)
where equality is achieved when the posterior distribution $P(G^{*}|X^{*})$ truly factors as in Eq. (58). Hence, when $M$ and $M^{*}$ are equal in distribution, $\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p}^{*})]\approx H(G^{*}|X^{*})$. Furthermore, the expected posterior loss is linearly related to the reconstructability, with a proportionality factor given by the entropy of $G^{*}$:
$\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p}^{*})]\approx H(G^{*})\Big{[}1-\Psi^{*}\Big{]}\,.$ (60)
Appendix E Bounds of the information gain
In this section, we show that the information gain is non-negative and bounded by the CE between the posterior and the prior. Recall that the information gain is given by Eq. (IV.1) (equivalently Eq. (30)):
$\mathcal{I}_{M}(x)=\mathbb{E}_{G|X=x}\quantity[\log\frac{P(G|X)}{P(G)}]\,.$